# Initial Guessing Bias: How Untrained Networks Favor Some Classes

### v4

@@ G.3.3 Amplification of IGB with depth
Here we prove that the distribution of $\mathbb{P}^{\mathcal{W}}\left(f_{0}\right)$ converges to a delta-peaked distribution in a multi-layer perceptron with ReLU activation when the number of layers goes to infinity. More precisely:
@@
Consider a dataset with Gaussian-distributed, i.i.d. components, where $\xi_{b}^{(a)}\sim\mathcal{N}(0,1)$, and $\xi_{b}^{(a)}$ represents the $b$-th component of the $a$-th input vector. Inputs are processed through an MLP, mapping the input to output according to Eq. (21), Eq. (22), and Eq. (23). The weights are initialized following the Kaiming normal scheme, i.e., $w^{(l)}_{ij}\sim\mathcal{N}\left(0,\frac{\sigma_{w}^{2}}{N_{l-1}}\right)$, with zero bias weights. Let us consider a multi-layer perceptron with $L$ hidden layers, ReLU activation function, and without pooling layers. Let us assume that, in the $\lim_{N_{l}\rightarrow\infty}\forall l\in[0,\dots,L]$, w.h.p.: $\textrm{Var}_{\chi}\left({h^{(l)}}\right)>0,\;\forall l\,$, (182) and in particular $\lim_{l\rightarrow\infty}\textrm{Var}_{\chi}\left({h^{(l)}}\right)>0,$ (183) where $\textrm{Var}_{\chi}\left({h^{(l)}}\right)$ indicate the variance of $\mathbb{P}^{\chi}\left(h^{(l)}\mid\mathcal{W}\right)$. Then $\lim_{L\rightarrow\infty}\gamma(\mathcal{A},\psi(\chi))=\infty\,.$ (184)
@@
$\textrm{Var}_{\mathcal{W}}\left({\left\langle{h^{(l+1)}}\right\rangle}\right)>(K+\epsilon)\textrm{Var}_{\mathcal{W}}\left({\left\langle{h^{(l)}}\right\rangle}\right)\,$, (185)
$\textrm{Var}_{\chi}\left({h^{(l+1)}}\right)<(K-\epsilon)\textrm{Var}_{\chi}\left({h^{(l)}}\right)\,$, (186)
@@
$\frac{\textrm{Var}_{\mathcal{W}}\left({\left\langle{h^{(l+2)}}\right\rangle}\right)}{\textrm{Var}_{\chi}\left({h^{(l+2)}}\right)}>\left(\frac{1}{K}\right)^{l}\frac{\textrm{Var}_{\mathcal{W}}\left({\left\langle{h^{(2)}}\right\rangle}\right)}{\textrm{Var}_{\chi}\left({h^{(2)}}\right)}\left(K+\epsilon\right)^{l}>l\underbrace{\left(\frac{1}{K}\frac{\textrm{Var}_{\mathcal{W}}\left({\left\langle{h^{(2)}}\right\rangle}\right)}{\textrm{Var}_{\chi}\left({h^{(2)}}\right)}\epsilon\right)}_{\equiv\frac{1}{C}}\,$,
$\Longrightarrow\frac{\textrm{Var}_{\chi}\left({h^{(l+2)}}\right)}{\textrm{Var}_{\mathcal{W}}\left({\left\langle{h^{(l+2)}}\right\rangle}\right)}<C\frac{1}{l}.$ (187)
@@
$\frac{\textrm{Var}_{\chi}\left({h^{(L)}}\right)}{\textrm{Var}_{\mathcal{W}}\left({\left\langle{h^{(L)}}\right\rangle}\right)}=\mathcal{O}\left(\frac{1}{L}\right)\Longrightarrow\frac{\textrm{Var}_{\mathcal{W}}\left({\mu_{c}}\right)}{\textrm{Var}_{\chi}\left({O^{(c)}}\right)}=\mathcal{O}\left(L\right)\Longrightarrow\lim_{L\rightarrow\infty}\frac{\textrm{Var}_{\mathcal{W}}\left({\mu_{c}}\right)}{\textrm{Var}_{\chi}\left({O^{(c)}}\right)}=\infty\,.$ (188)
@@
$\frac{1}{2}\textrm{Var}_{\chi}\left({h^{(l)}}\right)>\mkern 1.5mu\overline{\mkern-1.5mu\textrm{Var}_{\chi}\left({g^{(l)}}\right)\mkern-1.5mu}\mkern 1.5mu$ $\equiv\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{\left(g^{(l)}\right)^{2}}\right\rangle\mkern-1.5mu}\mkern 1.5mu-\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{g^{(l)}}\right\rangle^{2}\mkern-1.5mu}\mkern 1.5mu.$ (189)
To prove this inequality we will now focus separately on the two terms of the r.h.s.. First we analyze $\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{g^{(l)}}\right\rangle^{2}\mkern-1.5mu}\mkern 1.5mu$. From Eq. (176)
$\textrm{Var}_{\mathcal{W}}\left({\left\langle{h^{(l+1)}}\right\rangle}\right)=\sigma^{2}_{w}\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{g^{(l)}_{j}}\right\rangle^{2}\mkern-1.5mu}\mkern 1.5mu.$ (190)
$\begin{split}&\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{h^{(l)}_{j}}\right\rangle^{2}\mkern-1.5mu}\mkern 1.5mu=\int_{-\infty}^{\infty}p_{\left\langle{h^{(l)}_{j}}\right\rangle}\left({x}\right)x^{2}dx=\int_{-\infty}^{0}p_{\left\langle{h^{(l)}_{j}}\right\rangle}\left({x}\right)x^{2}dx+\int_{0}^{\infty}p_{\left\langle{h^{(l)}_{j}}\right\rangle}\left({x}\right)x^{2}dx\,.\end{split}$ (191)
We now use the symmetry of $p_{\left\langle{h^{(l)}_{j}}\right\rangle}\left({x}\right)$ (which is a Gaussian centered in 0), i.e.
$p_{\left\langle{h^{(l)}_{j}}\right\rangle}\left({x}\right)=p_{\left\langle{h^{(l)}_{j}}\right\rangle}\left({-x}\right)\,$, (192)
@@
$\begin{split}&\mu_{+}\left(\mu\right)=\int_{0}^{\infty}\mathcal{N}\left(x;\mu,\textrm{Var}_{\chi}\left({h^{(l+1)}}\right)\right)xdx>0\,,\\ &\mu_{-}\left(\mu\right)=\int_{-\infty}^{0}\mathcal{N}\left(x;\mu,\textrm{Var}_{\chi}\left({h^{(l+1)}}\right)\right)xdx<0\,.\end{split}$ (194)
@@
$\frac{1}{2}\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{h^{(l)}_{j}}\right\rangle^{2}\mkern-1.5mu}\mkern 1.5mu=\int_{0}^{\infty}p_{\left\langle{h^{(l)}_{j}}\right\rangle}\left({x}\right)\big(\mu_{+}^{2}+\mu_{-}^{2}+2\underbrace{\mu_{+}\mu_{-}}_{\leq 0}\big)dx.$ (195)
Note that $\mu_{+}\mu_{-}=0$ implies either $\mu_{+}=0$ or $\mu_{-}=0$ i.e. a distribution $p_{h^{(l)}_{j}}^{(\chi)}\left({x}\right)$ with non-positive or non-negative support. Since $p_{h^{(l)}_{j}}^{(\chi)}\left({x}\right)$ is a Gaussian distribution this may only happen in the limit of its variance going to 0 i.e. if the Gaussian shrink into a Dirac delta distribution. By hypothesis we excluded this possibility (Eq. (182) and Eq. (183)); therefore $\mu_{+}\mu_{-}<0$ and we can rewrite
$\frac{1}{2}\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{h^{(l)}_{j}}\right\rangle^{2}\mkern-1.5mu}\mkern 1.5mu<\underbrace{\int_{0}^{\infty}p_{\left\langle{h^{(l)}_{j}}\right\rangle}\left({x}\right)\left(\mu_{+}(x)^{2}+\mu_{-}(x)^{2}\right)dx}_{\equiv I}\;.$ (196)
Now let us focus on the integral $I$ of Eq. (196). We note that
$-\mu_{-}(x)$ $=-\int_{-\infty}^{0}\mathcal{N}\left(\tilde{x};x,\textrm{Var}_{\chi}\left({h^{(l+1)}}\right)\right)\tilde{x}d\tilde{x}==\int_{0}^{\infty}\mathcal{N}\left(y;-x,\textrm{Var}_{\chi}\left({h^{(l+1)}}\right)\right)ydy=\mu_{+}(-x)\,$, (197)
where in the second step we just changed the integration variable $\tilde{x}\rightarrow y\equiv-\tilde{x}$. We can therefore rewrite
@@
$\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{g^{(l)}_{j}}\right\rangle^{2}\mkern-1.5mu}\mkern 1.5mu>\left(\frac{1}{2}\right)\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{h^{(l)}_{j}}\right\rangle^{2}\mkern-1.5mu}\mkern 1.5mu\,.$ (199)
We now turn our attention to the second term appearing in the r.h.s. of Eq. (189), i.e. $\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{\left(g^{(l)}\right)^{2}}\right\rangle\mkern-1.5mu}\mkern 1.5mu$:
$\begin{split}\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{\left(g^{(l)}\right)^{2}}\right\rangle\mkern-1.5mu}\mkern 1.5mu&=\int_{-\infty}^{\infty}p_{\left\langle{h^{(l)}_{j}}\right\rangle}\left({x}\right)\left\langle{\left(g^{(l)}\right)^{2}}\right\rangle(x)\;dx=\int_{0}^{\infty}p_{\left\langle{h^{(l)}_{j}}\right\rangle}\left({x}\right)\left(\left\langle{\left(g^{(l)}\right)^{2}}\right\rangle(x)+\left\langle{\left(g^{(l)}\right)^{2}}\right\rangle(-x)\right)\;dx=\\ &=\int_{0}^{\infty}p_{\left\langle{h^{(l)}_{j}}\right\rangle}\left({x}\right)\left(\left\langle{\left(g^{(l)}\right)^{2}}\right\rangle_{+}(x)+\left\langle{\left(g^{(l)}\right)^{2}}\right\rangle_{-}(x)\right)\;dx,\end{split}$ (200)
@@
$\begin{split}&\left\langle{\left(g^{(l)}\right)^{2}}\right\rangle_{+}(x)=\int_{0}^{\infty}\mathcal{N}\left(y;x,\textrm{Var}_{\chi}\left({h^{(l+1)}}\right)\right)y^{2}dy\,,\\ &\left\langle{\left(g^{(l)}\right)^{2}}\right\rangle_{-}(x)=\int_{-\infty}^{0}\mathcal{N}\left(y;x,\textrm{Var}_{\chi}\left({h^{(l+1)}}\right)\right)y^{2}dy.\end{split}$ (201)
@@
$\left\langle{\left(g^{(l)}\right)^{2}}\right\rangle_{+}(x)+\left\langle{\left(g^{(l)}\right)^{2}}\right\rangle_{-}(x)=\left\langle{\left(h^{(l)}\right)^{2}}\right\rangle(x);$ (202)
@@
$\begin{split}&\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{\left(g^{(l)}_{j}\right)^{2}}\right\rangle\mkern-1.5mu}\mkern 1.5mu=\int_{0}^{\infty}p_{\left\langle{h^{(l)}_{j}}\right\rangle}\left({x}\right)\left\langle{\left(h^{(l)}_{j}\right)^{2}}\right\rangle(x)\;dx=\frac{1}{2}\int_{-\infty}^{\infty}p_{\left\langle{h^{(l)}_{j}}\right\rangle}\left({x}\right)\left\langle{\left(h^{(l)}_{j}\right)^{2}}\right\rangle(x)\;dx=\frac{1}{2}\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{\left(h^{(l)}_{j}\right)^{2}}\right\rangle\mkern-1.5mu}\mkern 1.5mu.\end{split}$ (203)
@@
$\begin{split}\mkern 1.5mu\overline{\mkern-1.5mu\textrm{Var}_{\chi}\left({g^{(l)}}\right)\mkern-1.5mu}\mkern 1.5mu&\equiv\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{\left(g^{(l)}\right)^{2}}\right\rangle\mkern-1.5mu}\mkern 1.5mu-\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{g^{(l)}}\right\rangle^{2}\mkern-1.5mu}\mkern 1.5mu=\frac{1}{2}\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{\left(h^{(l)}\right)^{2}}\right\rangle\mkern-1.5mu}\mkern 1.5mu-\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{g^{(l)}}\right\rangle^{2}\mkern-1.5mu}\mkern 1.5mu<\\ &<\frac{1}{2}\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{\left(h^{(l)}\right)^{2}}\right\rangle\mkern-1.5mu}\mkern 1.5mu-\frac{1}{2}\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{h^{(l)}}\right\rangle^{2}\mkern-1.5mu}\mkern 1.5mu=\frac{1}{2}\mkern 1.5mu\overline{\mkern-1.5mu\textrm{Var}_{\chi}\left({h^{(l)}}\right)\mkern-1.5mu}\mkern 1.5mu\xrightarrow{N_{l-1}\penalty 10000\ \rightarrow\penalty 10000\ \infty}\frac{1}{2}\textrm{Var}_{\chi}\left({h^{(l)}}\right).\end{split}$ (204)
where in the last step we used the concentration result discussed in App. C.2.1, i.e. that the distribution of $\textrm{Var}_{\chi}\left({h^{(l)}}\right)$ asymptotically narrows around its mean value, becoming, w.h.p. independent of the realization $\mathcal{W}$.
@@
$\textrm{Var}_{\chi}\left({h^{(l)}}\right)=\sigma_{w}^{2}\mkern 1.5mu\overline{\mkern-1.5mu\textrm{Var}_{\chi}\left({g^{(l-1)}_{j}}\right)\mkern-1.5mu}\mkern 1.5mu$. (205)
@@
$\textrm{Var}_{\chi}\left({h^{(l+1)}}\right)<\left(\frac{\sigma_{w}^{2}}{2}-\epsilon^{\prime}_{l}\right)\textrm{Var}_{\chi}\left({h^{(l)}}\right)\,$, (206)
@@
$\textrm{Var}_{\mathcal{W}}\left({\left\langle{h^{(l+1)}}\right\rangle}\right)>\left(\frac{\sigma_{w}^{2}}{2}+\epsilon_{l}\right)\textrm{Var}_{\mathcal{W}}\left({\left\langle{h^{(l)}}\right\rangle}\right)\,$, (207)
@@
$\begin{split}&\frac{\textrm{Var}_{\mathcal{W}}\left({\left\langle{h^{(l+2)}}\right\rangle}\right)}{\textrm{Var}_{\chi}\left({h^{(l+2)}}\right)}>\left(\frac{2}{\sigma_{w}^{2}}\right)^{l}\frac{\textrm{Var}_{\mathcal{W}}\left({\left\langle{h^{(2)}}\right\rangle}\right)}{\textrm{Var}_{\chi}\left({h^{(2)}}\right)}\left(\frac{\sigma_{w}^{2}}{2}+\epsilon\right)^{l}>l\underbrace{\left(\frac{2}{\sigma_{w}^{2}}\frac{\textrm{Var}_{\mathcal{W}}\left({\left\langle{h^{(2)}}\right\rangle}\right)}{\textrm{Var}_{\chi}\left({h^{(2)}}\right)}\epsilon\right)}_{\equiv\frac{1}{C}}\\ &\Longrightarrow\frac{\textrm{Var}_{\chi}\left({h^{(l+2)}}\right)}{\textrm{Var}_{\mathcal{W}}\left({\left\langle{h^{(l+2)}}\right\rangle}\right)}<C\frac{1}{l}.\end{split}$ (209)

### v5

@@ G.3.3 Amplification of IGB with depth
Here we prove that the distribution of $f_{G_{0}}\left({g_{0}}\right)$ converges to a delta-peaked distribution in a multi-layer perceptron with ReLU activation when the number of layers goes to infinity. More precisely:
@@
Consider a dataset with Gaussian-distributed, i.i.d. components, where $\xi_{b}^{(a)}\sim\mathcal{N}(0,1)$, and $\xi_{b}^{(a)}$ represents the $b$-th component of the $a$-th input vector. Inputs are processed through an MLP, mapping the input to output according to Eq. (21), Eq. (22), and Eq. (23). The weights are initialized following the Kaiming normal scheme, i.e., $w^{(l)}_{ij}\sim\mathcal{N}\left(0,\frac{\sigma_{w}^{2}}{N_{l-1}}\right)$, with zero bias weights. Let us consider a multi-layer perceptron with $L$ hidden layers, ReLU activation function, and without pooling layers. Let us assume that, in the $\lim_{N_{l}\rightarrow\infty}\forall l\in[0,\dots,L]$, w.h.p.: $\textrm{Var}_{\chi}\left({Z^{(l)}}\right)>0,\;\forall l\,$, (182) and in particular $\lim_{l\rightarrow\infty}\textrm{Var}_{\chi}\left({Z^{(l)}}\right)>0,$ (183) where $\textrm{Var}_{\chi}\left({Z^{(l)}}\right)$ indicate the variance of $f_{Z^{(l)}}^{(\chi)}\left({z{}{}}\right)$. Then $\lim_{L\rightarrow\infty}\gamma(\mathcal{A},\psi(\chi))=\infty\,.$ (184)
@@
$\textrm{Var}_{\mathcal{W}}\left({\left\langle{Z^{(l+1)}}\right\rangle}\right)>(K+\epsilon)\textrm{Var}_{\mathcal{W}}\left({\left\langle{Z^{(l)}}\right\rangle}\right)\,$, (185)
$\textrm{Var}_{\chi}\left({Z^{(l+1)}}\right)<(K-\epsilon)\textrm{Var}_{\chi}\left({Z^{(l)}}\right)\,$, (186)
@@
$\frac{\textrm{Var}_{\mathcal{W}}\left({\left\langle{Z^{(l+2)}}\right\rangle}\right)}{\textrm{Var}_{\chi}\left({Z^{(l+2)}}\right)}>\left(\frac{1}{K}\right)^{l}\frac{\textrm{Var}_{\mathcal{W}}\left({\left\langle{Z^{(2)}}\right\rangle}\right)}{\textrm{Var}_{\chi}\left({Z^{(2)}}\right)}\left(K+\epsilon\right)^{l}>l\underbrace{\left(\frac{1}{K}\frac{\textrm{Var}_{\mathcal{W}}\left({\left\langle{Z^{(2)}}\right\rangle}\right)}{\textrm{Var}_{\chi}\left({Z^{(2)}}\right)}\epsilon\right)}_{\equiv\frac{1}{C}}\,$,
$\Longrightarrow\frac{\textrm{Var}_{\chi}\left({Z^{(l+2)}}\right)}{\textrm{Var}_{\mathcal{W}}\left({\left\langle{Z^{(l+2)}}\right\rangle}\right)}<C\frac{1}{l}.$ (187)
@@
$\frac{\textrm{Var}_{\chi}\left({Z^{(L)}}\right)}{\textrm{Var}_{\mathcal{W}}\left({\left\langle{Z^{(L)}}\right\rangle}\right)}=\mathcal{O}\left(\frac{1}{L}\right)\Longrightarrow\frac{\textrm{Var}_{\mathcal{W}}\left({\mu_{c}}\right)}{\textrm{Var}_{\chi}\left({O_{c}}\right)}=\mathcal{O}\left(L\right)\Longrightarrow\lim_{L\rightarrow\infty}\frac{\textrm{Var}_{\mathcal{W}}\left({\mu_{c}}\right)}{\textrm{Var}_{\chi}\left({O_{c}}\right)}=\infty\,.$ (188)
@@
$\frac{1}{2}\textrm{Var}_{\chi}\left({Z^{(l)}}\right)>\mkern 1.5mu\overline{\mkern-1.5mu\textrm{Var}_{\chi}\left({H^{(l)}}\right)\mkern-1.5mu}\mkern 1.5mu$ $\equiv\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{\left(H^{(l)}\right)^{2}}\right\rangle\mkern-1.5mu}\mkern 1.5mu-\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{H^{(l)}}\right\rangle^{2}\mkern-1.5mu}\mkern 1.5mu.$ (189)
To prove this inequality we will now focus separately on the two terms of the r.h.s.. First we analyze $\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{H^{(l)}}\right\rangle^{2}\mkern-1.5mu}\mkern 1.5mu$. From Eq. (176)
$\textrm{Var}_{\mathcal{W}}\left({\left\langle{Z^{(l+1)}}\right\rangle}\right)=\sigma^{2}_{w}\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{H^{(l)}_{j}}\right\rangle^{2}\mkern-1.5mu}\mkern 1.5mu.$ (190)
$\begin{split}&\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{Z^{(l)}_{j}}\right\rangle^{2}\mkern-1.5mu}\mkern 1.5mu=\int_{-\infty}^{\infty}f_{\left\langle{Z^{(l)}_{j}}\right\rangle}\left({m_{z}}\right)m_{z}^{2}\,dm_{z}=\int_{-\infty}^{0}f_{\left\langle{Z^{(l)}_{j}}\right\rangle}\left({m_{z}}\right)m_{z}^{2}\,dm_{z}+\int_{0}^{\infty}f_{\left\langle{Z^{(l)}_{j}}\right\rangle}\left({m_{z}}\right)m_{z}^{2}\,dm_{z}\,.\end{split}$ (191)
We now use the symmetry of $f_{\left\langle{Z^{(l)}_{j}}\right\rangle}\left({m_{z}}\right)$ (which is a Gaussian centered in 0), i.e.
$f_{\left\langle{Z^{(l)}_{j}}\right\rangle}\left({m_{z}}\right)=f_{\left\langle{Z^{(l)}_{j}}\right\rangle}\left({-m_{z}}\right)\,$, (192)
@@
$\begin{split}&\mu_{+}\left(m_{z{}{}}\right)=\int_{0}^{\infty}\mathcal{N}\left(z{}{};m_{z{}{}},\textrm{Var}_{\chi}\left({Z^{(l+1)}}\right)\right)z{}{}dz{}{}>0\,,\\ &\mu_{-}\left(m_{z{}{}}\right)=\int_{-\infty}^{0}\mathcal{N}\left(z{}{};m_{z{}{}},\textrm{Var}_{\chi}\left({Z^{(l+1)}}\right)\right)z{}{}dz{}{}<0\,.\end{split}$ (194)
@@
$\frac{1}{2}\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{Z^{(l)}_{j}}\right\rangle^{2}\mkern-1.5mu}\mkern 1.5mu=\int_{0}^{\infty}f_{\left\langle{Z^{(l)}_{j}}\right\rangle}\left({x}\right)\big(\mu_{+}^{2}+\mu_{-}^{2}+2\underbrace{\mu_{+}\mu_{-}}_{\leq 0}\big)dx.$ (195)
Note that $\mu_{+}\mu_{-}=0$ implies either $\mu_{+}=0$ or $\mu_{-}=0$ i.e. a distribution $f_{Z^{(l)}_{j}}^{(\chi)}\left({z{}{}}\right)$ with non-positive or non-negative support. Since $f_{Z^{(l)}_{j}}^{(\chi)}\left({z{}{}}\right)$ is a Gaussian distribution this may only happen in the limit of its variance going to 0 i.e. if the Gaussian shrink into a Dirac delta distribution. By hypothesis we excluded this possibility (Eq. (182) and Eq. (183)); therefore $\mu_{+}\mu_{-}<0$ and we can rewrite
$\frac{1}{2}\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{Z^{(l)}_{j}}\right\rangle^{2}\mkern-1.5mu}\mkern 1.5mu<\underbrace{\int_{0}^{\infty}f_{\left\langle{Z^{(l)}_{j}}\right\rangle}\left({m_{z{}{}}}\right)\left(\mu_{+}(m_{z{}{}})^{2}+\mu_{-}(m_{z{}{}})^{2}\right)dm_{z{}{}}}_{\equiv I}\;.$ (196)
Now let us focus on the integral II of Eq. (196). We note that
$-\mu_{-}(m_{z{}{}})$ $=-\int_{-\infty}^{0}\mathcal{N}\left(z{}{};m_{z{}{}},\textrm{Var}_{\chi}\left({Z^{(l+1)}}\right)\right)z{}{}dz{}{}==\int_{0}^{\infty}\mathcal{N}\left(y;-m_{z{}{}},\textrm{Var}_{\chi}\left({Z^{(l+1)}}\right)\right)ydy=\mu_{+}(-m_{z{}{}})\,$, (197)
where in the second step we just changed the integration variable $z{}{}\rightarrow y\equiv-z{}{}$. We can therefore rewrite
@@
$\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{H^{(l)}_{j}}\right\rangle^{2}\mkern-1.5mu}\mkern 1.5mu>\left(\frac{1}{2}\right)\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{Z^{(l)}_{j}}\right\rangle^{2}\mkern-1.5mu}\mkern 1.5mu\,.$ (199)
We now turn our attention to the second term appearing in the r.h.s. of Eq. (189), i.e. $\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{\left(H^{(l)}\right)^{2}}\right\rangle\mkern-1.5mu}\mkern 1.5mu$:
$\begin{split}\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{\left(H^{(l)}\right)^{2}}\right\rangle\mkern-1.5mu}\mkern 1.5mu&=\int_{-\infty}^{\infty}f_{\left\langle{Z^{(l)}_{j}}\right\rangle}\left({m_{z{}{}}}\right)\left\langle{\left(H^{(l)}\right)^{2}}\right\rangle(m_{z{}{}})\;dm_{z{}{}}\\ &=\int_{0}^{\infty}f_{\left\langle{Z^{(l)}_{j}}\right\rangle}\left({m_{z{}{}}}\right)\left(\left\langle{\left(H^{(l)}\right)^{2}}\right\rangle(m_{z{}{}})+\left\langle{\left(H^{(l)}\right)^{2}}\right\rangle(-m_{z{}{}})\right)\;dm_{z{}{}}=\\ &=\int_{0}^{\infty}f_{\left\langle{Z^{(l)}_{j}}\right\rangle}\left({m_{z{}{}}}\right)\left(\left\langle{\left(H^{(l)}\right)^{2}}\right\rangle_{+}(m_{z{}{}})+\left\langle{\left(H^{(l)}\right)^{2}}\right\rangle_{-}(m_{z{}{}})\right)\;dm_{z{}{}},\end{split}$ (200)
@@
$\begin{split}&\left\langle{\left(H^{(l)}\right)^{2}}\right\rangle_{+}(m_{z{}{}})=\int_{0}^{\infty}\mathcal{N}\left(z{}{};m_{z{}{}},\textrm{Var}_{\chi}\left({Z^{(l+1)}}\right)\right)z{}{}^{2}dz{}{}\,,\\ &\left\langle{\left(H^{(l)}\right)^{2}}\right\rangle_{-}(m_{z{}{}})=\int_{-\infty}^{0}\mathcal{N}\left(z{}{};m_{z{}{}},\textrm{Var}_{\chi}\left({Z^{(l+1)}}\right)\right)z{}{}^{2}dz{}{}.\end{split}$ (201)
@@
$\left\langle{\left(H^{(l)}\right)^{2}}\right\rangle_{+}(m_{z{}{}})+\left\langle{\left(H^{(l)}\right)^{2}}\right\rangle_{-}(m_{z{}{}})=\left\langle{\left(Z^{(l)}\right)^{2}}\right\rangle(m_{z{}{}});$ (202)
@@
$\begin{split}&\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{\left(H^{(l)}_{j}\right)^{2}}\right\rangle\mkern-1.5mu}\mkern 1.5mu=\int_{0}^{\infty}f_{\left\langle{Z^{(l)}_{j}}\right\rangle}\left({m_{z{}{}}}\right)\left\langle{\left(Z^{(l)}_{j}\right)^{2}}\right\rangle(m_{z{}{}})\;dm_{z{}{}}=\frac{1}{2}\int_{-\infty}^{\infty}f_{\left\langle{Z^{(l)}_{j}}\right\rangle}\left({m_{z{}{}}}\right)\left\langle{\left(Z^{(l)}_{j}\right)^{2}}\right\rangle(m_{z{}{}})\;dm_{z{}{}}=\frac{1}{2}\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{\left(Z^{(l)}_{j}\right)^{2}}\right\rangle\mkern-1.5mu}\mkern 1.5mu.\end{split}$ (203)
@@
$\begin{split}\mkern 1.5mu\overline{\mkern-1.5mu\textrm{Var}_{\chi}\left({H^{(l)}}\right)\mkern-1.5mu}\mkern 1.5mu&\equiv\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{\left(H^{(l)}\right)^{2}}\right\rangle\mkern-1.5mu}\mkern 1.5mu-\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{H^{(l)}}\right\rangle^{2}\mkern-1.5mu}\mkern 1.5mu=\frac{1}{2}\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{\left(Z^{(l)}\right)^{2}}\right\rangle\mkern-1.5mu}\mkern 1.5mu-\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{H^{(l)}}\right\rangle^{2}\mkern-1.5mu}\mkern 1.5mu<\\ &<\frac{1}{2}\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{\left(Z^{(l)}\right)^{2}}\right\rangle\mkern-1.5mu}\mkern 1.5mu-\frac{1}{2}\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{Z^{(l)}}\right\rangle^{2}\mkern-1.5mu}\mkern 1.5mu=\frac{1}{2}\mkern 1.5mu\overline{\mkern-1.5mu\textrm{Var}_{\chi}\left({Z^{(l)}}\right)\mkern-1.5mu}\mkern 1.5mu\xrightarrow{N_{l-1}\penalty 10000\ \rightarrow\penalty 10000\ \infty}\frac{1}{2}\textrm{Var}_{\chi}\left({Z^{(l)}}\right).\end{split}$ (204)
where in the last step we used the concentration result discussed in App. C.2.1, i.e. that the distribution of $\textrm{Var}_{\chi}\left({Z^{(l)}}\right)$ asymptotically narrows around its mean value, becoming, w.h.p. independent of the realization $\mathcal{W}$.
@@
$\textrm{Var}_{\chi}\left({Z^{(l)}}\right)=\sigma_{w}^{2}\mkern 1.5mu\overline{\mkern-1.5mu\textrm{Var}_{\chi}\left({H^{(l-1)}_{j}}\right)\mkern-1.5mu}\mkern 1.5mu$. (205)
@@
$\textrm{Var}_{\chi}\left({Z^{(l+1)}}\right)<\left(\frac{\sigma_{w}^{2}}{2}-\epsilon^{\prime}_{l}\right)\textrm{Var}_{\chi}\left({Z^{(l)}}\right)\,$, (206)
@@
$\textrm{Var}_{\mathcal{W}}\left({\left\langle{Z^{(l+1)}}\right\rangle}\right)>\left(\frac{\sigma_{w}^{2}}{2}+\epsilon_{l}\right)\textrm{Var}_{\mathcal{W}}\left({\left\langle{Z^{(l)}}\right\rangle}\right)\,$, (207)
@@
$\begin{split}&\frac{\textrm{Var}_{\mathcal{W}}\left({\left\langle{Z^{(l+2)}}\right\rangle}\right)}{\textrm{Var}_{\chi}\left({Z^{(l+2)}}\right)}>\left(\frac{2}{\sigma_{w}^{2}}\right)^{l}\frac{\textrm{Var}_{\mathcal{W}}\left({\left\langle{Z^{(2)}}\right\rangle}\right)}{\textrm{Var}_{\chi}\left({Z^{(2)}}\right)}\left(\frac{\sigma_{w}^{2}}{2}+\epsilon\right)^{l}>l\underbrace{\left(\frac{2}{\sigma_{w}^{2}}\frac{\textrm{Var}_{\mathcal{W}}\left({\left\langle{Z^{(2)}}\right\rangle}\right)}{\textrm{Var}_{\chi}\left({Z^{(2)}}\right)}\epsilon\right)}_{\equiv\frac{1}{C}}\\ &\Longrightarrow\frac{\textrm{Var}_{\chi}\left({Z^{(l+2)}}\right)}{\textrm{Var}_{\mathcal{W}}\left({\left\langle{Z^{(l+2)}}\right\rangle}\right)}<C\frac{1}{l}.\end{split}$ (209)